one end of a glass prism is in the shape of a triangle with a height that is 1 inch longer than twice the base. if the area of the triangle is 39 square inches, then how long are the base and height

Answer by Boreal(15235)

A=1/2 bh
but h=2b+1
Therefore, 39=(1/2)(2b^2+b)
or 2b^2+b=78
2b^2+b-78=0
(2b+13)(b-6)=0
b=6 as only positive root
h=13